A ball is tossed off the side of a building. The ball is given an initial horizontal velocity of 14.5 m/s and the ground. How many meters from the base of the building will the ball land?


1
Expert's answer
2021-09-15T11:38:56-0400
"x=vt\\x=\sqrt{\frac{2h}{g}}v\\x=\sqrt{\frac{2(44)}{9.8}}14.5=43.5\ m"